# Balanced Space- and Time-based Duty-cycle Scheduling for Light-based IoT

### v1

@@ 3.3. BST-TDMA algorithm
As a general conventional approach to scheduling the next duty-cycle of energy-harvesting LIoT nodes, we used U-STDMA scheduling algorithm, which schedules a cluster of the LIoT nodes by adding the $T_{s}$ + $dur$ (duty-cycle duration) and checks if the $next_{time}$ has a conflict with other nodes’ duty-cycles. When a conflict is detected the $next_{time}$ is calculated by adding the difference between $dur$ and $g$, Alg.3.
Algorithm 3 U-STDMA
  while True do
     updated $\leftarrow$ false
     for $i=1$ to length$(node_{C})$ do
@@
        $last_{time}$leftarrow result(2, result(1,:) == $event_{id}$)
[…]
        while $next_{time}$leq $end_{time}$ do
           for i = 1 to length(result) do
              $scheduled_{time}$leftarrow result(2, i)
              if abs($scheduled_{time}$ $next_{time}$) $>$ $dur$ then
                 result $\leftarrow$ [result, [$event_{id}$; $next_{time}$]]
                 updated $\leftarrow$ true
                 break
              else
                 $conflicting_{time}$leftarrow result(2, abs(result(2,:) -$next_{time}$) $<$ $dur$)
                 $g$leftarrow $\min$(abs($conflicting_{time}$ $next_{time}$))
                 $next_{time}$leftarrow $next_{time}$ + ($dur$ $g$)
              end if
           end for
        end while
     end for
     if not updated then
        break
     end if
  end while
4. Performance Evaluation
In this section, we present the performance evaluation of the BST-TDMA method. We compare it with the unbalanced scheduling approach U-STDMA, which maximizes the number of duty-cycles. We consider that the duty-cycles of a cluster of LIoT nodes are scheduled by the gateway node per each AP using Alg.1. By using the whitepaper of EHU manufacturer (epishine, ), we can calculate the EHU harvested power $P_{harv}$ based on the linear relationship between the $P_{harv}$ and illuminance. Using the measurements of the LIoT node power profiling, we measured $E_{dev}$ = 0.2233 J during duty-cycle lasting $dur$ = 4.45 seconds. Also, our LIoT nodes consume $P_{sleep}$ = 0.28mW power during sleep time. Based on our measurements, we can calculate $T_{s}$ as (2). A visualization of the LIoT nodes’ transmissions over time when using BST-TDMA is presented in Fig. 3, in which one can see the cluster $node_{C}$ of 5 LIoT nodes with $T_{s}=[306,235,666,505,546]$ and $per=47$ can be seen. Fig.3 illustrates every LIoT node schedule for a duration of 30mins and intervals between duty-cycles of the same node are not equal.
Figure 3. LIoT nodes scheduling of a single cluster.
Figure 4. Combined LIoT schedule.
[…]
Figure 5. BST-STDMA vs U-STDMA CDF.
Figure 6. Duty-cycle intervals mode comparison by number of nodes.
Figure 7. Duty-cycle intervals mode comparison by hours.
A comparison of both scheduling approaches with a large number of LIoT nodes shows, that the mode of the scheduled duty-cycle intervals for BST-TDMA is always equal to calculated $per$, while the U-STDMA shows randomness in this term, Fig.6.
Another significant difference between the two MAC protocols is found while testing them with increasing measurement time from 1 to 60 hours. Fig.7 shows, that the mode interval between duty-cycles is always equal to $per$ in the case of BST-TDMA, and for the U-STDMA, it increases more than 5 times over $per$ after 9 hours for the same cluster of LIoT $node_{C}$.
5. Conclusion
In this paper, we introduced the BST-TDMA duty-cycle scheduling algorithm for intermittent operating LIoT nodes, which optimizes the duty-cycle schedule by taking into account the spatial distribution of LIoT nodes to enhance network efficiency, reduce communication blind spots, and ensure comprehensive environmental sensing. The simulation results demonstrate the effectiveness of the BST-TDMA algorithm compared to the conventional Unbalanced scheduling approach. Our approach offers a sustainable and scalable solution for large-scale indoor IoT applications and significantly reduces the randomness of duty-cycle interval duration. Future work will focus on further refining the BST-TDMA algorithm to adapt to dynamic environmental conditions and varying node densities. Additionally, we plan to explore the duty-cycle scheduling based on the type of environment observation.
6. Acknowledgments
The present work has received funding from the European Union’s Horizon 2020 Marie Skłodowska Curie Innovative Training Network Greenedge (GA. No. 953775). The work of Hazem Sallouha was funded by the Research Foundation – Flanders (FWO), Postdoctoral Fellowship No. 12ZE [^2] N.

### v2

@@ III-C BST-TDMA algorithm
IV Performance Evaluation
In this section, we present the performance evaluation of the BST-TDMA method. We compare it with the unbalanced scheduling approach presented in [9], which maximizes the number of duty-cycles without knowing measurement distribution over space-time. We assume that the duty-cycles of a set of LIoT nodes per cluster are scheduled by the gateway node using Alg.1. A visualization of the LIoT nodes’ transmissions over time when using BST-TDMA is presented in Fig. 3, in which one can see the set $node_{C}$ of 5 LIoT nodes with $T_{s}=[306,235,666,505,546]$ and $per=47$ can be seen. Fig.3 illustrates every LIoT node schedule for a duration of 30mins and intervals between duty-cycles of the same node are not equal.
Figure 3: LIoT nodes scheduling of a single cluster.
Figure 4: Combined LIoT schedule.
@@ IV Performance Evaluation
Figure 5: BST-STDMA vs U-STDMA CDF
A comparison of both scheduling approaches with a large number of LIoT nodes shows, that the mode of the scheduled duty-cycle intervals for BST-TDMA is always equal to calculated $per$, while the U-STDMA shows randomness in this term, Fig.6. Another significant difference between the two MAC protocols is found while testing them with increasing measurement time from 1 to 60 hours. Fig.7 shows, that the mode interval between duty-cycles is always equal to $per$ in the case of BST-TDMA, and for the U-STDMA, it increases more than 5 times over $per$ after 9 hours for the same set of LIoT $node_{C}$.
Figure 6: duty-cycle interval mode comparison by number of nodes.
Figure 7: duty-cycle interval mode comparison by duration.
V Conclusion
In this paper, we introduced the BST-TDMA duty-cycle scheduling algorithm for intermittent operating LIoT nodes, which optimizes the duty-cycle schedule taking into account the spatial distribution of LIoT nodes to enhance network efficiency, reduce communication blind spots, and ensure comprehensive environmental sensing. The simulation results demonstrate the effectiveness of the BST-TDMA algorithm compared to the conventional Unbalanced scheduling approach. Our approach offers a sustainable and scalable solution for large-scale indoor IoT applications, addressing the challenges associated with batteryless LIoT nodes and VLC-based communication networks. Future work will focus on further refining the BST-TDMA algorithm to adapt to dynamic environmental conditions and varying node densities. Additionally, we plan to explore the duty-cycle scheduling based on the type of environment observation.